# Orbifold modifications of complex analytic varieties

### v4

@@ 6.A. Preparatory lemmas
$W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\}\to Z_{\operatorname{\textsubscript{\rm sm}}}.$
Moreover, we can choose $W$ so that the natural morphism $\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})$ is isomorphic.
Let $\eta\colon Z\to Y\times\check{S}^{\circ}$ be the morphism defined by $\varphi\circ\psi^{-1}$, which is an open embedding. Then $\eta$ can be written in coordinates as
$\eta\colon(x,y,z,s^{\prime})\mapsto(x^{\prime},y^{\prime},z^{\prime},s^{\prime}).$
@@
$\eta_{*}\colon\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})\to H_{2}$
@@
$\varphi_{*}=(\eta\circ\psi)_{*}\colon\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
induces an isomorphism from $\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})$ to $H_{2}$. This implies that the morphism $\varphi_{*}\colon\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}}\times S^{\circ})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ induces an isomorphism from $H_{1}$ to $H_{2}$.
Since $u$ is connected to $\mathbf{0}_{3}$ by path in $U$, we see that $V$ is deformation equivalent to $\{\mathbf{0}_{3}\}\times S^{\circ}$ inside $U\times S^{\circ}$. Thus $\varphi(V)$ is deformation equivalent to $\{\mathbf{0}_{3}\}\times S^{\circ}$ inside $Y\times\check{S}^{\circ}$. Hence by Lemma 6.2, if $q\colon Y\times\check{S}^{\circ}\to\check{S}^{\circ}$ is the natural projection, then $q\circ\varphi$ induces an isomorphism from $\pi_{1}(V)$ to the image of $\pi_{1}(S^{\circ})$ in $\pi_{1}(\check{S}^{\circ})$. This implies that $\varphi_{*}$ induces an injective map on $G_{1}$. Since $G_{1}$ is a free Abelian group, we deduce that the intersection $\varphi_{*}(G_{1})\cap\varphi_{*}(H_{1})$ is trivial. Hence $\varphi_{*}$ is injective. Finally, we notice that the image of $\varphi_{*}$ is equal to the preiamge of $\pi_{1}(S^{\circ})$ under the morphism
$q_{*}\colon\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})\to\pi_{1}(\check{S}^{\circ}),$
@@
On the other hand, the quotient map $E\to E/E^{\prime}$ admits a splitting $E/E^{\prime}\to E$. Let $J\subseteq E$ be its image. Then $E=JE^{\prime}$ and $J\cap E^{\prime}=\{e\}$, where $e$ is the neutral element of $E$. Let $K^{\prime}$ be a complement of $H$ in $E^{\prime}$. Then $K:=JK^{\prime}$ is a complement of $H$ in $E$. Furthermore, $K\cap E^{\prime}=K^{\prime}$. It follows that $N\cap E^{\prime}\subseteq N^{\prime}$. This completes the proof of the lemma. ∎
@@ 6.B. Construction of covering spaces
    where $l$ is an integer, $\alpha(s)$, $\beta(s)$ and $\gamma(s)$ are holomorphic functions on $\check{S}$, whose zero loci are contained in $H$, and $u(y,s)$ is a holomorphic function defined on $\mathbb{D}\times\check{S}$. The composite inclusion $\check{S}\subseteq X\subseteq\mathbb{D}^{3}\times\check{S}$ is identified with $\{\mathbf{0}_{3}\}\times\check{S}\subseteq\mathbb{D}^{3}\times\check{S}$. Furthermore, if $\theta\colon M\to\check{S}$ is the projection sending $(x,y,z,s)$ to $s$, then $\Delta=\theta^{-1}(H)\cap X$. Let $\check{S}^{\circ}=\check{S}\setminus H$, $M^{\circ}=\theta^{-1}(\check{S}^{\circ})$, and $X^{\circ}=\theta^{-1}(\check{S}^{\circ})\cap X=X\setminus\Delta$.
@@
    Let $o_{S}\in\check{S}$ be any point. We will work in a neighborhood of $o_{S}$ in $X$. Up to shrinking $X$ around $o_{S}$, up to replacing $G$ by the stabilizer of $o_{S}$, and up to reducing the number $k$, we assume that $\check{S}$ an open neighborhood of the origin $\mathbf{0}_{n}=o_{S}$ of $\mathbb{C}^{n}$ contained in a polydisc of radius smaller than 1. Since $\check{S}$ is smooth, the action of $G$ on the germ $(o_{S}\in\check{S})$ is isomorphic to a linear action (see [Car54, Lemme 1] or [Kaw88, Lemma 9.9]). Therefore, we may assume that $\check{S}$ is stable under the scaling with any positive factor less than 1. In particular, it is contractible. We can also assume that there is a coordinates system $(T_{1},...,T_{n})$ on $\check{S}$, such that $H_{i}=\{T_{i}=0\}$, with $i=1,...,k$. Let $T=T_{1}\cdots T_{k}$.
@@
    Let $F^{\prime}(x,y,z)$ be the polynomial function obtained by replacing $\alpha,\beta,\gamma$ with 1 and replacing $u$ with 0 in the expression of $F$. We denote by $Y$ the hypersurface in $\mathbb{C}^{3}$ defined by $F^{\prime}=0$. Then $Y$ has an isolated Du Val singularity at $\mathbf{0}_{3}$. Let $U\subseteq Y$ be a bounded open neighborhood of $\mathbf{0}_{3}$, such that the natural morphism of fundamental groups $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$ is isomorphic.
The open subset $M^{\prime}$ in the item (4) of the previous setup plays an important role in our proof of Proposition 6.7. The reason that we work with two neighborhood $S$ and $\check{S}$ is following one. The projection $\theta\colon X\to\check{S}$ is not $G$-invariant in general. As a consequence, $X$ may not contain an open subset of the shape $((\mathbb{D}^{\prime})^{3}\times\check{S})\cap M$.
@@
With the notation of Setup 6.6, there is an integer $N\geq 0$, such that the following properties hold for any integer $N^{\prime}\geq N$. Let $\rho\colon\widehat{X}\to X$ be the blowup of $X$ at the $H_{i}$’s as in Subsection 3.C, such that there are $6N^{\prime}$ blowups whose centers are $H_{i}$ for all $i$. Particularly, there is a natural action of $G$ on $\widehat{X}$ so that $\rho$ is $G$-equivariant. Let $\widehat{S}$ be the strict transform of $S$ in $\widehat{X}$, let $\widehat{H}\subseteq\widehat{S}$ be the preimage of $H$, and let $\widehat{X}^{\circ}=\rho^{-1}(X^{\circ})=\widehat{X}\setminus\rho^{-1}(\Delta)$.
Then there is an open neighborhood $W$ of $\widehat{S}$ and a finite cover $\nu\colon\overline{W}\to W$, which satisfy the following properties.
@@
    The finite morphism $\nu$ is functorial in the following sense. For any point $\hat{o}\in W$, let $o=\rho(\hat{o})\in X$. Assume that there is an open neighborhood $W^{\prime}$ of $\hat{o}$ in $W$, such that there is a finite morphism $\nu^{\prime}\colon\overline{W}^{\prime}\to W^{\prime}$, which is constructed by the same method on a neighborhood $X^{\prime}$ of $\rho(W^{\prime})$ in $X$. Then there is an open neighborhood $W^{\prime\prime}$ of $\hat{o}$ in $W^{\prime}$, such that every connected component of $\nu^{-1}(W^{\prime\prime})$ is isomorphic to every connected component $(\nu^{\prime})^{-1}(W^{\prime\prime})$.
@@
We will apply this local construction on orbifold charts of some complex analytic orbispaces. The item (5) above is to ensure that the constructions are compatible along the overlaps. To explain the assumption on the finite morphism $\nu^{\prime}\colon\overline{W}^{\prime}\to W^{\prime}$ in this item, we first illustrate the construction of the finite morphism $\nu\colon\overline{W}\to W$. Initially, we construct an open embedding $X^{\circ}\to Y\times\check{S}^{\circ}$, such that its restriction on $\check{S}^{\circ}$ is the identity map onto $\{\mathbf{0}_{3}\}\times\check{S}^{\circ}$. We note that $\check{S}\subseteq\mathbb{C}^{n}$, and $\check{S}^{\circ}\subseteq\mathbb{C}^{n}\setminus H$, where we still denote by $H$ its Zariski closure in $\mathbb{C}^{n}$, which is again the union of some coordinate hyperplanes. Let $Z\to Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ}$ be the étale finite morphism induced by the canonical subgroup in Lemma 6.5 of the fundamental group $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$. Then we construct $\overline{W}\to W$ as the unique finite morphism induced by the fiber product $W^{\circ}_{\operatorname{\textsubscript{\rm sm}}}\times_{X^{\circ}}Z$, see Theorem 2.4. Now for the morphism $\nu\colon\overline{W}^{\prime}\to W^{\prime}$ in the item (5), the assumption particularly implies that there is an open embedding $X^{\prime}\cap X^{\circ}\to Y\times(\mathbb{C}^{n}\setminus H^{\prime})$, where $H^{\prime}$ is the union of some coordinate hyperplanes, such that $X^{\prime}\cap S^{\circ}$ is mapped to $\{\mathbf{0}_{3}\}\times(\mathbb{C}^{n}\setminus H^{\prime})$. The morphism $\nu^{\prime}$ is induced by taking the fiber product of $W^{\prime}_{\operatorname{\textsubscript{\rm sm}}}\cap W^{\circ}$ over the canonical finite morphism $Z^{\prime}\to Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{\prime})$ as above.
@@
We will divide the proof of the proposition into a sequence of lemmas. The sketch is as follows. As explained in the previous paragraph, we first construct an open embedding $X^{\circ}\to Y\times\check{S}^{\circ}$. We may construct a finite morphism $\overline{X}\to X$ by taking basechange as in the previous paragraph. However, since we have little information on the fundamental group of $X^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$, it is not clear to us that $\overline{X}$ satisfies the item (4) and (5) of the proposition, which signify that $\overline{X}$ is independent of the open embedding $X^{\circ}\to Y\times\check{S}^{\circ}$. To deal with this problem, we will construct explicitly a tubular neighborhood $V$ of $S^{\circ}$ in $X^{\circ}$, such that $\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})\cong\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})\times\mathbb{Z}^{k}\cong\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$. Then the finite cover $\overline{V}\to V$ constructed by this method is independent of the choice of the open embedding from $X^{\circ}$ to $Y\times\check{S}^{\circ}$. A new issue is that, $V$ may not extend to an open neighborhood of $S$ in $X$. It can collapse when approaching to $H\subseteq S$. Thus $\overline{V}$ may not extend to a covering space over any neighborhood of $S$. Our solution to this is to blow up $X$ at the $H_{i}$’s as in Subsection 3.C. We show that, after some blowups, the preimage of $V$ extends to some open neighborhood $W$ of the strict transform of $S$. Then the covering space $\overline{V}$ extends to a finite cover $\overline{W}\to W$.
@@
$\varphi\colon(\overline{x},\overline{y},\overline{z},s)\mapsto(a(s)\overline{x},b(s)\overline{y},\zeta(\overline{y},\overline{z},s),s)=(x,y,z,s),$
@@
$\varphi^{-1}\colon(x,y,z,s)\mapsto(a(s)^{-1}x,b(s)^{-1}y,c(s)^{-1}(\gamma(s)z+u(y,s)y),s)=(\overline{x},\overline{y},\overline{z},s).$
@@
$\eta\colon(\overline{x},\overline{y},\overline{z},s)\mapsto(T^{p}\overline{x},T^{q}\overline{y},T^{r}\overline{z},s),$
@@
$\eta\colon(\overline{x},\overline{y},\overline{z},s)\mapsto(T^{p}\overline{x},T^{q}\overline{y},T^{r}\overline{z},s),$
@@
With the notation above, there exists an integer $N>0$ such that the following property holds. Let $\rho\colon\widehat{M}\to M$ be a morphism obtained by blowing up $M$ at the $H_{i}^{\prime}$s as in Subsection 3.C, such that there are at least $N$ blowups whose centers are $H_{i}$ for all $i$. Let $\widehat{X}=\rho^{-1}(X)$, $\widehat{X}^{\circ}=\rho^{-1}(X^{\circ})$, $\widehat{V}=\rho^{-1}(V)$ and let $\widehat{S}$ be the strict transform of $S$ in $\widehat{M}$. Then there is an open neighborhoods $W$ of $\widehat{S}$ in $\widehat{X}$, such that $W^{\circ}:=W\cap\widehat{X}^{\circ}$ is contained in $\widehat{V}$, and that $W$ is invariant under the natural action of $G$ on $\widehat{X}$. Moreover, we have $\rho(W\cap\rho^{-1}(\Delta))\subseteq H\cap S$.
@@
There is some $\varepsilon>0$ such that the intersection $Y\cap(\mathbb{D}_{\varepsilon})^{3}$ is contained in $U$. Then $\varphi((Y\cap(\mathbb{D}_{\varepsilon})^{3})\times S^{\circ})\subseteq V$. The application $\varphi^{-1}\colon M^{\circ}\to Y\times S^{\circ}$ can be written with coordinates as
$\varphi^{-1}\colon(x,y,z,s)\mapsto(a(s)^{-1}x,b(s)^{-1}y,\xi(y,z,s),s),$
@@
are holomorphic functions on $\check{S}$. Hence they are bounded on $S$. Therefore there is some $\delta>0$ such that if $\widehat{M}^{\circ}:=\rho^{-1}(M^{\circ})$ and if $\hat{o}=(\hat{x},\hat{y},\hat{z},s)$ is contained in $\widehat{M}^{\circ}\cap((\mathbb{D}_{\delta})^{3}\times\widehat{S})$, then
@@
Assume that $\hat{o}\in\widehat{W}\cap\rho^{-1}(\Delta)=\widehat{W}\setminus\widehat{M}^{\circ}$ is a point with coordinate $(\hat{x},\hat{y},\hat{z},s)$ in $(\mathbb{D}_{\delta})^{3}\times\widehat{S}$. Then $\hat{T}(s)=0$ and $\rho(\hat{o})$ has coordinates $(0,0,0,s)$ in $M$. This implies that $\rho(\hat{o})\in H\cap S\subseteq X$. In particular, we see that $\widehat{W}\subseteq\widehat{X}$. Hence we can define $W=\bigcap_{g\in G}g(\widehat{W})$. Then $W$ is a $G$-invariant neighborhood of $\widehat{S}$ in $\widehat{X}$. This completes the proof of the lemma. ∎
@@
With the notation in Lemma 6.10, we fix such a bimeromorphic morphism $\rho\colon\widehat{X}\to X$ such that there are $6N^{\prime}$ blowups whose centers are $H_{i}$ for all $i$, where $N^{\prime}\geq N$ is an integer. Then there is a sequence of morphisms of fundamental groups
@@
We denote by $\iota\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ the composition of the previous sequence. Then $\iota$ is surjective.
Furthermore, if $\psi\colon X^{\circ}\to Y\times\check{S}^{\circ}$ is an open embedding, whose restriction on $\check{S}^{\circ}\subseteq X^{\circ}$ is equal to the identity map from $\check{S}^{\circ}$ to $\{\mathbf{0}_{3}\}\times\check{S}^{\circ}$. Then the morphism of fundamental groups
$(\psi\circ\rho)_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$

### v6

@@ 6.A. Preparatory lemmas
$W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\}\rightarrow Z_{\operatorname{\textsubscript{\rm sm}}}.$
Moreover, we can choose $W$ so that the natural morphism $\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})$ is isomorphic.
Let $\eta\colon Z\rightarrow Y\times\check{S}^{\circ}$ be the morphism defined by $\varphi\circ\psi^{-1}$, which is an open embedding. Then $\eta$ can be written in coordinates as
$\eta\colon(x,y,z,s^{\prime})\mapstochar\rightarrow(x^{\prime},y^{\prime},z^{\prime},s^{\prime}).$
@@
$\eta_{*}\colon\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})\rightarrow H_{2}$
@@
$\varphi_{*}=(\eta\circ\psi)_{*}\colon\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$
induces an isomorphism from $\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})$ to $H_{2}$. This implies that the morphism $\varphi_{*}\colon\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}}\times S^{\circ})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ induces an isomorphism from $H_{1}$ to $H_{2}$.
Since $u$ is connected to $\mathbf{0}_{3}$ by path in $U$, we see that $V$ is deformation equivalent to $\{\mathbf{0}_{3}\}\times S^{\circ}$ inside $U\times S^{\circ}$. Thus $\varphi(V)$ is deformation equivalent to $\{\mathbf{0}_{3}\}\times S^{\circ}$ inside $Y\times\check{S}^{\circ}$. Hence by Lemma 6.2, if $q\colon Y\times\check{S}^{\circ}\rightarrow\check{S}^{\circ}$ is the natural projection, then $q\circ\varphi$ induces an isomorphism from $\pi_{1}(V)$ to the image of $\pi_{1}(S^{\circ})$ in $\pi_{1}(\check{S}^{\circ})$. This implies that $\varphi_{*}$ induces an injective map on $G_{1}$. Since $G_{1}$ is a free Abelian group, we deduce that the intersection $\varphi_{*}(G_{1})\cap\varphi_{*}(H_{1})$ is trivial. Hence $\varphi_{*}$ is injective. Finally, we notice that the image of $\varphi_{*}$ is equal to the preiamge of $\pi_{1}(S^{\circ})$ under the morphism
$q_{*}\colon\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})\rightarrow\pi_{1}(\check{S}^{\circ}),$
@@
On the other hand, the quotient map $E\rightarrow E/E^{\prime}$ admits a splitting $E/E^{\prime}\rightarrow E$. Let $J\subseteq E$ be its image. Then $E=JE^{\prime}$ and $J\cap E^{\prime}=\{e\}$, where $e$ is the neutral element of $E$. Let $K^{\prime}$ be a complement of $H$ in $E^{\prime}$. Then $K:=JK^{\prime}$ is a complement of $H$ in $E$. Furthermore, $K\cap E^{\prime}=K^{\prime}$. It follows that $N\cap E^{\prime}\subseteq N^{\prime}$. This completes the proof of the lemma. ∎
@@ 6.B. Construction of covering spaces
    where $l$ is an integer, $\alpha(s)$, $\beta(s)$ and $\gamma(s)$ are holomorphic functions on $\check{S}$, whose zero loci are contained in $H$, and $u(y,s)$ is a holomorphic function defined on $\mathbb{D}\times\check{S}$. The composite inclusion $\check{S}\subseteq X\subseteq\mathbb{D}^{3}\times\check{S}$ is identified with $\{\mathbf{0}_{3}\}\times\check{S}\subseteq\mathbb{D}^{3}\times\check{S}$. Furthermore, if $\theta\colon M\rightarrow\check{S}$ is the projection sending $(x,y,z,s)$ to $s$, then $\Delta=\theta^{-1}(H)\cap X$. Let $\check{S}^{\circ}=\check{S}\setminus H$, $M^{\circ}=\theta^{-1}(\check{S}^{\circ})$, and $X^{\circ}=\theta^{-1}(\check{S}^{\circ})\cap X=X\setminus\Delta$.
@@
    Let $o_{S}\in\check{S}$ be any point. We will work in a neighborhood of $o_{S}$ in $X$. Up to shrinking $X$ around $o_{S}$, up to replacing $G$ by the stabilizer of $o_{S}$, and up to reducing the number $k$, we assume that $\check{S}$ an open neighborhood of the origin $\mathbf{0}_{n}=o_{S}$ of $\mathbb{C}^{n}$ contained in a polydisc of radius smaller than 1. Since $\check{S}$ is smooth, the action of $G$ on the germ $(o_{S}\in\check{S})$ is isomorphic to a linear action (see [Car54, Lemme 1] or [Kaw88, Lemma 9.9]). Therefore, we may assume that $\check{S}$ is stable under the scaling with any positive factor less than 1. In particular, it is contractible. We can also assume that there is a coordinates system $(T_{1},...,T_{n})$ on $\check{S}$, such that $H_{i}=\{T_{i}=0\}$, with i=1,,ki=1,...,k. Let $T=T_{1}\cdots T_{k}$.
@@
    Let $F^{\prime}(x,y,z)$ be the polynomial function obtained by replacing $\alpha,\beta,\gamma$ with 1 and replacing $u$ with 0 in the expression of $F$. We denote by $Y$ the hypersurface in $\mathbb{C}^{3}$ defined by $F^{\prime}=0$. Then $Y$ has an isolated Du Val singularity at $\mathbf{0}_{3}$. Let $U\subseteq Y$ be a bounded open neighborhood of $\mathbf{0}_{3}$, such that the natural morphism of fundamental groups $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$ is isomorphic.
The open subset $M^{\prime}$ in the item (4) of the previous setup plays an important role in our proof of Proposition 6.7. The reason that we work with two neighborhood $S$ and $\check{S}$ is following one. The projection $\theta\colon X\rightarrow\check{S}$ is not $G$-invariant in general. As a consequence, $X$ may not contain an open subset of the shape $((\mathbb{D}^{\prime})^{3}\times\check{S})\cap M$.
@@
With the notation of Setup 6.6, there is an integer $N\geq 0$, such that the following properties hold for any integer $N^{\prime}\geq N$. Let $\rho\colon\widehat{X}\rightarrow X$ be the blowup of $X$ at the $H_{i}$’s as in Subsection 3.C, such that there are $6N^{\prime}$ blowups whose centers are $H_{i}$ for all $i$. Particularly, there is a natural action of $G$ on $\widehat{X}$ so that $\rho$ is $G$-equivariant. Let $\widehat{S}$ be the strict transform of $S$ in $\widehat{X}$, let $\widehat{H}\subseteq\widehat{S}$ be the preimage of $H$, and let $\widehat{X}^{\circ}=\rho^{-1}(X^{\circ})=\widehat{X}\setminus\rho^{-1}(\Delta)$.
Then there is an open neighborhood $W$ of $\widehat{S}$ and a finite cover $\nu\colon\overline{W}\rightarrow W$, which satisfy the following properties.
@@
    The finite morphism $\nu$ is functorial in the following sense. For any point $\hat{o}\in W$, let $o=\rho(\hat{o})\in X$. Assume that there is an open neighborhood $W^{\prime}$ of $\hat{o}$ in $W$, such that there is a finite morphism $\nu^{\prime}\colon\overline{W}^{\prime}\rightarrow W^{\prime}$, which is constructed by the same method on a neighborhood $X^{\prime}$ of $\rho(W^{\prime})$ in $X$. Then there is an open neighborhood $W^{\prime\prime}$ of $\hat{o}$ in $W^{\prime}$, such that every connected component of $\nu^{-1}(W^{\prime\prime})$ is isomorphic to every connected component $(\nu^{\prime})^{-1}(W^{\prime\prime})$.
@@
We will apply this local construction on orbifold charts of some complex analytic orbispaces. The item (5) above is to ensure that the constructions are compatible along the overlaps. To explain the assumption on the finite morphism $\nu^{\prime}\colon\overline{W}^{\prime}\rightarrow W^{\prime}$ in this item, we first illustrate the construction of the finite morphism $\nu\colon\overline{W}\rightarrow W$. Initially, we construct an open embedding $X^{\circ}\rightarrow Y\times\check{S}^{\circ}$, such that its restriction on $\check{S}^{\circ}$ is the identity map onto $\{\mathbf{0}_{3}\}\times\check{S}^{\circ}$. We note that $\check{S}\subseteq\mathbb{C}^{n}$, and $\check{S}^{\circ}\subseteq\mathbb{C}^{n}\setminus H$, where we still denote by $H$ its Zariski closure in $\mathbb{C}^{n}$, which is again the union of some coordinate hyperplanes. Let $Z\rightarrow Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ}$ be the étale finite morphism induced by the canonical subgroup in Lemma 6.5 of the fundamental group $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$. Then we construct $\overline{W}\rightarrow W$ as the unique finite morphism induced by the fiber product $W^{\circ}_{\operatorname{\textsubscript{\rm sm}}}\times_{X^{\circ}}Z$, see Theorem 2.4. Now for the morphism $\nu\colon\overline{W}^{\prime}\rightarrow W^{\prime}$ in the item (5), the assumption particularly implies that there is an open embedding $X^{\prime}\cap X^{\circ}\rightarrow Y\times(\mathbb{C}^{n}\setminus H^{\prime})$, where $H^{\prime}$ is the union of some coordinate hyperplanes, such that $X^{\prime}\cap S^{\circ}$ is mapped to $\{\mathbf{0}_{3}\}\times(\mathbb{C}^{n}\setminus H^{\prime})$. The morphism $\nu^{\prime}$ is induced by taking the fiber product of $W^{\prime}_{\operatorname{\textsubscript{\rm sm}}}\cap W^{\circ}$ over the canonical finite morphism $Z^{\prime}\rightarrow Y_{\operatorname{\textsubscript{\rm sm}}}\times(\mathbb{C}^{n}\setminus H^{\prime})$ as above.
@@
We will divide the proof of the proposition into a sequence of lemmas. The sketch is as follows. As explained in the previous paragraph, we first construct an open embedding $X^{\circ}\rightarrow Y\times\check{S}^{\circ}$. We may construct a finite morphism $\overline{X}\rightarrow X$ by taking basechange as in the previous paragraph. However, since we have little information on the fundamental group of $X^{\circ}_{\operatorname{\textsubscript{\rm sm}}}$, it is not clear to us that $\overline{X}$ satisfies the item (4) and (5) of the proposition, which signify that $\overline{X}$ is independent of the open embedding $X^{\circ}\rightarrow Y\times\check{S}^{\circ}$. To deal with this problem, we will construct explicitly a tubular neighborhood $V$ of $S^{\circ}$ in $X^{\circ}$, such that $\pi_{1}(V_{\operatorname{\textsubscript{\rm sm}}})\cong\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})\times\mathbb{Z}^{k}\cong\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$. Then the finite cover $\overline{V}\rightarrow V$ constructed by this method is independent of the choice of the open embedding from $X^{\circ}$ to $Y\times\check{S}^{\circ}$. A new issue is that, $V$ may not extend to an open neighborhood of $S$ in $X$. It can collapse when approaching to $H\subseteq S$. Thus $\overline{V}$ may not extend to a covering space over any neighborhood of $S$. Our solution to this is to blow up $X$ at the $H_{i}$’s as in Subsection 3.C. We show that, after some blowups, the preimage of $V$ extends to some open neighborhood $W$ of the strict transform of $S$. Then the covering space $\overline{V}$ extends to a finite cover $\overline{W}\rightarrow W$.
@@
$\varphi\colon(\overline{x},\overline{y},\overline{z},s)\mapstochar\rightarrow(a(s)\overline{x},b(s)\overline{y},\zeta(\overline{y},\overline{z},s),s)=(x,y,z,s),$
@@
$\varphi^{-1}\colon(x,y,z,s)\mapstochar\rightarrow(a(s)^{-1}x,b(s)^{-1}y,c(s)^{-1}(\gamma(s)z+u(y,s)y),s)=(\overline{x},\overline{y},\overline{z},s).$
@@
$\eta\colon(\overline{x},\overline{y},\overline{z},s)\mapstochar\rightarrow(T^{p}\overline{x},T^{q}\overline{y},T^{r}\overline{z},s),$
@@
$\eta\colon(\overline{x},\overline{y},\overline{z},s)\mapstochar\rightarrow(T^{p}\overline{x},T^{q}\overline{y},T^{r}\overline{z},s),$
@@
With the notation above, there exists an integer $N>0$ such that the following property holds. Let $\rho\colon\widehat{M}\rightarrow M$ be a morphism obtained by blowing up $M$ at the $H_{i}^{\prime}$s as in Subsection 3.C, such that there are at least $N$ blowups whose centers are $H_{i}$ for all $i$. Let $\widehat{X}=\rho^{-1}(X)$, $\widehat{X}^{\circ}=\rho^{-1}(X^{\circ})$, $\widehat{V}=\rho^{-1}(V)$ and let $\widehat{S}$ be the strict transform of $S$ in $\widehat{M}$. Then there is an open neighborhoods $W$ of $\widehat{S}$ in $\widehat{X}$, such that $W^{\circ}:=W\cap\widehat{X}^{\circ}$ is contained in $\widehat{V}$, and that $W$ is invariant under the natural action of $G$ on $\widehat{X}$. Moreover, we have $\rho(W\cap\rho^{-1}(\Delta))\subseteq H\cap S$.
@@
There is some $\varepsilon>0$ such that the intersection $Y\cap(\mathbb{D}_{\varepsilon})^{3}$ is contained in $U$. Then $\varphi((Y\cap(\mathbb{D}_{\varepsilon})^{3})\times S^{\circ})\subseteq V$. The application $\varphi^{-1}\colon M^{\circ}\rightarrow Y\times S^{\circ}$ can be written with coordinates as
$\varphi^{-1}\colon(x,y,z,s)\mapstochar\rightarrow(a(s)^{-1}x,b(s)^{-1}y,\xi(y,z,s),s),$
@@
[…]
@@
[…]
@@
With the notation in Lemma 6.10, we fix such a bimeromorphic morphism $\rho\colon\widehat{X}\rightarrow X$ such that there are $6N^{\prime}$ blowups whose centers are $H_{i}$ for all $i$, where $N^{\prime}\geq N$ is an integer. Then there is a sequence of morphisms of fundamental groups
@@
We denote by $\iota\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ the composition of the previous sequence. Then $\iota$ is surjective.
Furthermore, if $\psi\colon X^{\circ}\rightarrow Y\times\check{S}^{\circ}$ is an open embedding, whose restriction on $\check{S}^{\circ}\subseteq X^{\circ}$ is equal to the identity map from $\check{S}^{\circ}$ to $\{\mathbf{0}_{3}\}\times\check{S}^{\circ}$. Then the morphism of fundamental groups
$(\psi\circ\rho)_{*}\colon\pi_{1}(W^{\circ}_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$